# EdgeOL: Efficient in-situ Online Learning on Edge Devices

### 30 Jan 2024 v1

@@ 3.3. Intra-tuning
A Preliminary Exploration. We conduct a preliminary experiment to show the effect of layer freezing on execution time, energy consumption, and accuracy. As an example, we employ ResNet50 and MobileNetV2 on the “ImgNet to CF10” test case and follow the experimental setup in Section 3.1, and we use the fine-tuning strategy $S7$.
@@
Figure 7 shows the accuracy when freezing different numbers of layers of ResNet50 and MobileNetV2. For a more comprehensive exploration, we randomly selected two groups of layers to freeze, denoted as Config 1 and Config 2, respectively. First, we observe that as more layers are frozen, the accuracy drops accordingly. However, when we compare the accuracy of the two configurations, even with the same number of frozen layers, there is still a significant difference in accuracy. The most intriguing observation is that, when freezing layers appropriately (e.g., freezing 10 layers), the accuracy is not decreased and, in fact, increases over the non-freezing baseline. This is because freezing layers can i) avoid over-adaptation to training data and ii) improve model convergence speed (details in Figure 14 and Section 5.1). This indicates that layer freezing is not a simple trade-off between accuracy and efficiency. As the preliminary experimental results, we use relatively simple settings to explore the feasibility and potential of incorporating layer freezing in online learning. Note that similar trends are observed with other experimental settings (e.g., using other fine-tuning strategies).
Even though layer freezing has promising performance in time and energy efficiency and accuracy in online learning, several critical questions remain. Firstly, to incorporate layer freezing in the fine-tuning process, determining which layers are appropriate to freeze is a challenging problem. Secondly, there are also many choices about when to freeze a layer.
@@ 4. EdgeOL Design
acc_im_this_ft
Validation acc. improvement in this fine-tuning round
acc_im_last_ft
Validation acc. improvement in last fine-tuning round
acc_im_next_ft
Approx. validation acc. improvement in next fine-tuning round
@@
Based on these insights, we propose EdgeOL, an efficient online learning framework for edge devices. Figure 8 shows the overview of the EdgeOL framework to achieve energy efficiency and high inference accuracy through i) inter-tuning and ii) intra-tuning optimizations. Specifically, for inter-tuning, we propose a novel Dynamic and Adaptive Fine-tuning Frequency (DAF) design that dynamically and adaptively adjusts the fine-tuning frequency to reduce the execution time and energy consumption (Section 4.1). For intra-tuning, we propose a similarity-guided freezing (SimFreeze) method to automatically freeze layers during online learning to save computation and memory costs while improving accuracy (Section 4.2). If there exist multiple changes in the deployment scenario during online learning, EdgeOL can also detect them and quickly adapt to the new scenarios (Section 4.3). Moreover, it can utilize unlabeled data via semi-supervised learning (Section 4.4). The EdgeOL optimization design is described in Algorithm 1 with terminology and abbreviations listed in Table 2.
Figure 8. Overview of EdgeOL. ①, ②, and ③ in Figure 8b indicate the occurrence of freezing, matching the case 1, 2, and 3 in Figure 2, respectively. ④ indicates the occurrence of unfreezing right after scenario change is detected.
@@ 4.1.1. Key Design Factors
Cost-Effectiveness. As discussed in Section 3.2 and Figure 5, each time launching a fine-tuning round will inevitably introduce extra time and energy overheads. Therefore, we must consider the cost-effectiveness. It indicates whether the potential model accuracy improvement of launching a fine-tuning round is worth the incurred overheads.
@@ 4.1.2. DAF Design Principle
To consider both cost-effectiveness and inference intensity, the proposed DAF dynamically orchestrates the fine-tuning frequency based on training data availability, the trend of the model’s validation accuracy, and the intensity of inference requests. It is important to emphasize that validation accuracy differs from inference accuracy of inference requests. Validation accuracy is obtained by evaluating model on validation dataset, which is formed by randomly separating a small portion ($\sim$5%) of the streaming-in training data while maintaining the same data distribution (ponce2006dataset, ). We cannot use inference accuracy because, in real-world applications, the inference requests will not have the corresponding ground truth labels; thus, we use validation accuracy to indicate model performance.
@@
Tuning down the fine-tuning frequency. In general, as the model gradually converges through multiple fine-tuning rounds during online learning, the cost-effectiveness of the fine-tuning rounds decreases (see Section 3.2.3). Therefore, the DAF design monitors the improvements in the validation accuracy of each fine-tuning round. If the improvement is insufficient to remain cost-effective, the fine-tuning frequency is tuned down accordingly (lines 12 to 17 in Algorithm 1).
Tuning up the fine-tuning frequency. To effectively improve inference accuracy under an intensive inference period, we tune up the fine-tuning frequency in DAF via a popular logarithmic-based function (manaseer2009logarithmic, ). It is calculated as $d=d*(1-1/log(d))$, where $d$ represents the number of data batches needed to trigger a fine-tuning. If inference requests are intensive, the $batches\_needed$ will be decreased, thereby tuning up the fine-tuning frequency (lines 20 to 23 in Algorithm 1). We opt for a logarithmic-based function to decrease $batches\_needed$ because it provides a moderate adjustment compared to two other prevalent value-adjusting functions: exponential-based function (kwak2005performance, ) and additive-based function (dumas2002markovian, ). The logarithmic-based function is less aggressive than the exponential-based function, yet more aggressive than the additive-based function.
Figure 9. An example of adaptive adjustment using DAF.
@@ 4.1.3. Case study
Figure 9 shows a real example of how DAF adaptively adjusts the fine-tuning frequency. The results are obtained on ResNet50 with a “ImgNet to CF10” task change scenario. From the figure, we have the following observations: ① shows that $batches\_needed$ has small values and remains unchanged for several consecutive fine-tuning rounds. This is because significant accuracy improvements are achieved at the beginning of the online learning process, and our DAF infers to keep a high fine-tuning frequency. ② shows obvious decreases in $batches\_needed$ as the DAF responds to the intensive incoming inference requests at those moments. Moreover, DAF increases the fine-tuning frequency to keep the model up-to-date. ③ shows the overall increasing trend of $batches\_needed$ throughout the online learning process since the model has generally converged, and DAF decreases the fine-tuning frequency (increases $batches\_needed$) to facilitate higher cost-effectiveness and reduce energy consumption.
@@ 10            $freeze\_interval\leftarrow freeze\_interval\times(1-1/log(freeze\_interval))$ ;
13             if $acc\_im\_this\_ft\leq 0$ $\mathrm{or}$ $acc\_im\_last\_ft\leq 0$ then
14                   $batches\_needed\leftarrow batches\_needed\times 2$;
@@
16            else if $acc\_im\_this\_ft\leq acc\_im\_last\_ft$ then
17                   $acc\_im\_next\_ft\leftarrow acc\_im\_this\_ft\times(acc\_im\_this\_ft/acc\_im\_last\_ft)$;
18                   $batches\_needed\leftarrow batches\_needed\times(acc\_im\_this\_ft/acc\_im\_next\_ft)$;
19
20
21
22# Inference # Tuning up the fine-tuning frequency if inference arrives then
@@ 25             $batches\_needed\leftarrow batches\_needed\times(1-1/log(batches\_needed))$ ;
26
27
28# Multiple scenario changes if a scenario change is detected then
@@ 4.2. Similarity-Guided Freezing (SimFreeze)
We next design SimFreeze that adaptively freezes appropriate layers during online learning.
Utilizing self-representational similarity to guide freezing. SimFreeze uses a layer’s self-representational similarity to guide whether it can be frozen. We consider the initial model before fine-tuning as the reference model. We define the self-representational similarity of a layer as the degree of similarity between the output feature maps of a layer in the current model version and the output feature maps of that layer in the reference model. As fine-tuning proceeds, the layers of the model are updated over time, and their self-representational similarity is also recorded. When a layer’s self-representational similarity is stabilized, then we consider the layer to have converged and it can be frozen.
To measure the self-representational similarity of two layers from two models, we use a recently proposed metric Centered Kernel Alignment (CKA) (kornblith2019similarity, ). The CKA value is obtained by comparing the output feature maps of two layers using the same input image batch. It can be calculated as:
@@
Since calculating CKA requires using the same input data, we collect the first arrived training data batch as our CKA test data. The CKA test data will be used as the input for the models to generate output feature maps of each layer. As shown in Algorithm 1 (lines 5 to 7), periodically (e.g., every 300 iterations), we calculate the CKA and check the self-representational similarity for each active (non-frozen) layer, which is the first step of the freezing process in Figure 8b. For each active layer, we compare the CKA value of the current model to the CKA value calculated last time. The layers whose CKA variations are below the threshold (e.g., 1%) are considered converged and will be frozen (lines 8 and 9 in Algorithm 1), as the second freezing step in Figure 8b.
@@
Figure 10 shows an example of the trend of CKA values of layers 1, 10, 15, 40, and 50 as fine-tuning proceeds. It can be observed that different layers require a different number of training iterations to converge. For example, layer 1 converges at the very beginning, while layer 50 continues to fluctuate at the end of the fine-tuning. Moreover, it is interesting to observe that later layers can converge faster than earlier layers (e.g., layer 15 vs. layer 10). This is due to residual connections in the model network architecture, making some later layers behave like earlier layers (veit2016residual, ). These observations show the feasibility and necessity of freezing layers in an adaptive manner rather than sequentially from front to back. Besides, we can also observe that once a layer has converged, its CKA value will remain stable and will not fluctuate significantly again. Therefore, if a layer is frozen, then it would be good to keep it frozen for higher energy efficiency unless there are changes in the model deployment scenario (discussed in Section 4.3).
Figure 10. CKA variation curve as fine-tuning proceeds. Result is obtained by fine-tuning ResNet50 on “ImgNet to CF10” test case.
Dynamic interval to conduct freezing processes. SimFreeze continuously conducts freezing processes (i.e., tracks the CKA variation and freezes the converged layers). However, conducting the freezing process too frequently (e.g., every training iteration) can lead to excessive overhead, as CKA calculation requires time and energy. Therefore, we opt for periodically tracking the CKA variation of active layers and freezing the converging ones. Further, we propose using a logarithmic-based function that progressively decreases the interval before subsequent freezing process after each freezing process (line 10 in Algorithm 1). This approach takes into consideration that more and more layers converge as fine-tuning proceeds. For example, layers 10 and 40 in Figure 10 have not converged at early training iterations but gradually converged. The log-based function is calculated as $n=n*(1-1/log(n))$, where $n$ represents the interval to conduct a freezing process.
4.3. Handling Multiple Scenario Changes
It is possible that the model undergoes multiple deployment scenario changes during online learning. It is crucial for the model to be updated quickly to adapt to these changes to deliver satisfactory results. In general, the scenario changes can be identified using many different methods. For example, one can track the distribution difference of the data in streaming-in inference requests using methods such as Least-Squares Density Difference (LSDD) (gama2014survey, ) or Maximum Mean Discrepancy (MMD) (gretton2012kernel, ). Or it can also be detected by a stand-alone sensor module in a comprehensive system (e.g., autonomous vehicle system (gormer2009vision, )). EdgeOL is compatible with these detection methods, and it is not our focus in this work.
@@ 4.3. Handling Multiple Scenario Changes
Once the scenario change is detected, at the inter-tuning level, we reset the fine-tuning frequency to the initial default value, thereby increasing the fine-tuning frequency. At the intra-tuning level, it may not be necessary to unfreeze all the frozen layers since some front layers are task-agnostic (discussed in Section 3.3). To decide which layers to unfreeze, we first update the CKA test data with the new scenario data (line 27 in Algorithm 1). Then, for each frozen layer, we compute its CKA using both original and new scenario CKA test data. If the CKA variation of a layer exceeds the threshold after the scenario change, it indicates that the feature extraction ability of that layer for new scenario data and previous scenario data is significantly different. In this case, we unfreeze that unstable layer to allow it to adapt to the new scenario (lines 28 to 31 in Algorithm 1).

### 16 Mar 2024 v2

@@ 3.2. Intra-tuning
A Preliminary Exploration. We conduct a preliminary experiment to show the effect of layer freezing on execution time, energy consumption, and accuracy. As an example, we employ ResNet50 and MobileNetV2 on the NC benchmark and follow the experimental setup in Section 5.1, and we use the strategy $S5$.
@@
Figure 7 shows the accuracy when freezing different numbers of layers of ResNet50 and MobileNetV2. For a more comprehensive exploration, we randomly selected two groups of layers to freeze, denoted as Config 1 and Config 2, respectively. First, we observe that as more layers are frozen, the accuracy drops accordingly. However, when we compare the accuracy of the two configurations, even with the same number of frozen layers, there is still a significant difference in accuracy. The most intriguing observation is that, when freezing layers appropriately (e.g., freezing 10 layers), the accuracy is not decreased and, in fact, increases over the non-freezing baseline. This is because freezing layers can i) avoid over-adaptation to training data and ii) improve model convergence speed (details in Figure 14 and Section 5.2). This indicates that layer freezing is not a simple trade-off between accuracy and efficiency. As the preliminary experimental results, we use relatively simple settings to explore the feasibility and potential of incorporating layer freezing in online learning. Note that similar trends are observed with other experimental settings (e.g., using other online learning strategies with different fine-tuning frequencies).
Even though layer freezing has promising performance in time and energy efficiency and accuracy in online learning, several critical questions remain. Firstly, to incorporate layer freezing in the fine-tuning process, determining which layers are appropriate to freeze is a challenging problem. Secondly, there are also many choices about when to freeze a layer. In addition, when the scenario changes, we also need to decide whether to resume training on certain frozen layers to quickly adapt to the new scenario.
@@ 4. EdgeOL Design
Based on the insights, we propose EdgeOL, an efficient online learning framework for edge devices. Figure 8 shows the overview of the EdgeOL framework to achieve energy efficiency and high inference accuracy through i) inter-tuning and ii) intra-tuning optimizations. Specifically, for inter-tuning, we propose a novel Dynamic and Adaptive Fine-tuning Frequency (DAF) design that dynamically and adaptively adjusts the fine-tuning frequency to reduce the execution time and energy consumption (Section 4.1). For intra-tuning, we propose a similarity-guided freezing (SimFreeze) method to automatically freeze/unfreeze layers during online learning to save computation costs while improving accuracy (Section 4.2). Moreover, EdgeOL can also use unlabeled data through semi-supervised learning techniques to enhance model performance without the need for extensive labeled data (Section 4.3). The EdgeOL optimization design is described in Algorithm 1 with terminology and abbreviations listed in Table 2.
Figure 8. Overview of EdgeOL. ①, ②, and ③ in Figure 8b indicate the occurrence of freezing, matching the case 1, 2, and 3 in Figure 2, respectively. ④ indicates the occurrence of unfreezing right after a scenario change.
@@ 4.1.1. Key Design Factors
Cost-Effectiveness. As discussed in Section 3.1, each time launching a fine-tuning round will inevitably introduce extra time and energy overheads. Therefore, we must consider the cost-effectiveness. It indicates whether the potential model accuracy improvement of launching a fine-tuning round is worth the incurred overheads.
@@
Scenario Changes. The model may undergo multiple deployment scenario changes during online learning. It is crucial for the model to update quickly to adapt to these changes to deliver satisfactory results.
In general, the scenario changes can be identified by many different methods. For example, one can track the distribution difference of the data in streaming-in inference requests using methods such as Least-Squares Density Difference (LSDD) (gama2014survey, ), Maximum Mean Discrepancy (MMD) (gretton2012kernel, ), and many other out-of-distribution data detection approaches (wu2022energy, ; liu2020energy, ; hendrycks2018deep, ; chen2021atom, ; hsu2020generalized, ). Or it can also be detected by a stand-alone sensor module in a comprehensive system (e.g., robotics system (lesort2020continual, ; zhao2021general, )). EdgeOL is compatible with these detection methods, and it is not our focus in this work.
@@ 4.1.2. DAF Design Principle
To take the above-mentioned design factors into account, the proposed DAF dynamically orchestrates the fine-tuning frequency based on training data availability, the trend of the model’s validation accuracy, the intensity of inference requests, and the changes in the model deployment scenario. It is important to emphasize that validation accuracy differs from inference accuracy of inference requests. Validation accuracy is obtained by evaluating the model on a validation dataset, which is formed by randomly separating a small portion ($\sim$5%) of the streaming-in training data while maintaining the same data distribution (ponce2006dataset, ). We cannot use inference accuracy because, in real-world applications, the inference requests will not have the corresponding ground truth labels; thus, we use validation accuracy to indicate model performance.
@@
Tuning down the frequency considering cost-effec- tiveness. In general, within one scenario, as the model gradually converges through multiple fine-tuning rounds during online learning, the cost-effectiveness of the fine-tuning rounds decreases (see Section 4). Therefore, the fine-tuning frequency should be gradually tuned down to remain cost-effective. Specifically, since launching each fine-tuning round incurs similar overhead, we consider maintaining cost-effectiveness as letting each round achieve comparable accuracy improvements. Accordingly, after a fine-tuning round, DAF estimates the amount of training data needed for the next round to achieve similar accuracy gains as the most recent round (lines 12 and 13 in Algorithm 1).
Inspired by prior works (bhardwaj2022ekya, ; peng2018optimus, ; mahajan2020themis, ), this estimation is achieved by employing a logistic regression model (hosmer2013applied, ), which extrapolate the accuracy improvement curve from data collected from prior fine-tuning rounds, as described by:
(1) $A(t)=\frac{L}{1+e^{-k(t-t_{0})}}$
where $A(t)$ is the estimated accuracy at training iteration $t$, $L$ represents the curve’s asymptote representing the maximum achievable accuracy, $k$ is the average accuracy growth rate, $t_{0}$ is the inflection point where the accuracy improvement rate begins to decline, and $e$ is the base of the natural logarithm. DAF employs the Non-Negative Least Squares (NNLS) solver (scipy.optimize.nnls, ) to fit this accuracy curve to the (training iteration, validation accuracy) data points collected from previous fine-tuning rounds. Each data point encapsulates the number of training iterations the model has experienced and the achieved validation accuracy at that iteration.
This model is then used to predict future accuracy improvements. This enables DAF to ensure sufficient accuracy gains in subsequent fine-tuning rounds, thereby maintaining cost-effectiveness. Typically, this is achieved by strategically increasing $batch\_needed$ (i.e., tuning down the fine-tuning frequency). Note that the accuracy might drop after a fine-tuning round due to the fluctuating nature of model training. In this case, DAF will use the accuracy gains in the previous round as the accuracy improvement target for the next round.
Tuning up the frequency considering inference intensity. To effectively improve inference accuracy under an intensive inference period, we tune up the fine-tuning frequency in DAF via a popular logarithmic-based function (manaseer2009logarithmic, ). It is calculated as $d=d*(1-1/log(d))$, where $d$ represents the number of data batches needed to trigger a fine-tuning round. If inference requests are intensive, the $batches\_needed$ will be quickly decreased, thereby tuning up the fine-tuning frequency (lines 16 to 19 in Algorithm 1). We opt for a logarithmic-based function to decrease $batches\_needed$ because it provides a moderate adjustment compared to two other prevalent value-adjusting functions: exponential-based function (kwak2005performance, ) and additive-based function (dumas2002markovian, ). The logarithmic-based function is less aggressive than the exponential-based function, yet more aggressive than the additive-based function.
Resetting the frequency upon a scenario change. Once a scenario change is detected, to ensure a quick adaptation to the new scenario to deliver satisfactory results, we increase the model fine-tuning frequency by resetting the frequency to the initial value.
Figure 9. An example of adaptive adjustment using DAF. The result is obtained by fine-tuning ResNet50 on the NC benchmark of the CORe50 dataset. The model is well-trained on the first scenario and then experiences subsequent scenarios. Here we show the 2nd and 3rd scenarios as an example. The red dotted line indicates a scenario change.
@@ 4.1.3. Case study
Figure 9 shows a real example of how DAF adaptively adjusts the fine-tuning frequency in two consecutive scenarios. From the figure, we have the following observations: ① shows that $batches\_needed$ remains a small value for several fine-tuning rounds. This is because significant accuracy improvements are achieved at the beginning of the online learning process in each scenario, and our DAF intends to keep a high fine-tuning frequency. ② shows obvious decreases in $batches\_needed$ as the DAF responds to the intensive incoming inference requests at those moments. Moreover, DAF increases the fine-tuning frequency to keep the model up-to-date. ③ shows the overall increasing trend of $batches\_needed$ throughout the online learning process in each scenario since the model has generally converged, and DAF decreases the fine-tuning frequency (increases $batches\_needed$) to facilitate higher cost-effectiveness and reduce energy consumption. ④ shows a significant decrease in $batches\_needed$ upon a scenario change, as DAF increases the fine-tuning frequency by setting it to the initial value. This ensures quick model adaption in the new scenario.
@@ 10            $freeze\_interval\leftarrow freeze\_interval\times(1-1/log(freeze\_interval))$ ;
13             $batch\_needed$leftarrow batch_needed_estimation();
14
@@ 13             $batch\_needed$leftarrow batch_needed_estimation();
16# Inference # Tuning up the fine-tuning frequency if inference arrives then
@@ 19             $batches\_needed\leftarrow batches\_needed\times(1-1/log(batches\_needed))$ ;
20
21
22# Handling scenario changes if a scenario change is detected then
@@ 4.2. Similarity-Guided Freezing (SimFreeze)
We next design SimFreeze that adaptively freezes and unfreezes appropriate layers during online learning. In general, as the model gradually converges as training proceeds within one scenario, SimFreeze identifies and freezes those converged layers. Upon encountering a scenario change, SimFreeze selectively resumes training on previously frozen layers that become unstable in the new scenario, facilitating a rapid and efficient adaptation to the changes.
4.2.1. Utilizing self-representational similarity to guide layer freezing.
SimFreeze uses a layer’s self-representational similarity to guide whether it can be frozen. We consider the initial model before fine-tuning as the reference model. We define the self-representational similarity of a layer as the degree of similarity between the output feature maps of a layer in the current model version and the output feature maps of that layer in the reference model. As fine-tuning proceeds, the layers of the model are updated over time, and their self-representational similarity is also recorded. When a layer’s self-representational similarity is stabilized, then we consider the layer to have converged and it can be frozen.
To measure the self-representational similarity of two layers from two models, we use a widely-used metric Centered Kernel Alignment (CKA) (kornblith2019similarity, ). The CKA value is obtained by comparing the output feature maps of two layers using the same input image batch. It can be calculated as:
@@
Instead of comparing to the initial model, an intuitive alternative is to use models from prior training iterations as reference models. In this case, we can compare the current model to the model in earlier training iterations to monitor the CKA variation trend. However, this requires regular updates of the reference model as fine-tuning proceeds, leading to an increase in memory writes and consequently higher energy consumption. As such, we would not adopt this method in our framework.
@@
Within a scenario, we collect the first arrived training data batch as the CKA test data for that scenario. The CKA test data will be used as the input for the models to generate output feature maps of each layer. As shown in Algorithm 1 (lines 5 to 7), periodically (e.g., every 200 iterations), we calculate the CKA and check the self-representational similarity for each active (non-frozen) layer, which is the first step of the freezing process in Figure 8b. For each active layer, we compare the CKA value of the current model to the CKA value calculated last time. The layers whose CKA variations are below the threshold (e.g., 1%) are considered converged and will be frozen (lines 8 and 9 in Algorithm 1), as the second freezing step in Figure 8b.
Figure 10. CKA variation curve as fine-tuning proceeds. The result is obtained by fine-tuning ResNet50 on the NC benchmark. The model is well-trained on the first scenario and then experiences subsequent scenarios. Here we show the CKA value in the 2nd and 3rd scenarios as an example.
@@
Figure 10 shows an example of the trend of CKA values of layers 1, 10, 15, 40, and 50 as fine-tuning proceeds. From the figure, we have the following observations. First, different layers require a different number of training iterations to converge. For example, layer 1 converges at the very beginning, while layer 50 fluctuates all the time. Moreover, it is interesting to observe that later layers can converge faster than earlier layers (e.g., layer 15 vs. layer 10). This is due to residual connections in the model network architecture, making some later layers behave like earlier layers (veit2016residual, ). These observations show the feasibility and necessity of freezing layers in an adaptive manner rather than sequentially from front to back. Besides, we can also observe that once a layer has converged, its CKA value will remain stable and will not fluctuate significantly again within the same scenario. Therefore, if a layer is frozen, then it would be good to keep it frozen for higher energy efficiency unless there are changes in the model deployment scenario.
4.2.2. Unfreezing layers upon scenario changes.
@@
Once a scenario change is detected, we need to resume training on certain frozen layers to ensure a quick adaptation. However, it may not be necessary to unfreeze all the frozen layers since some front layers are task-agnostic (discussed in Section 3.2). To decide which layers to unfreeze, we first update the CKA test data with the new scenario data (line 23 in Algorithm 1). Then, for each frozen layer, we compute its CKA using both original and new scenario CKA test data (lines 24 to 25 in Algorithm 1), as the first step of the unfreezing process in Figure 8b. If the CKA variation of a layer exceeds the threshold after the scenario change, it indicates that the feature extraction ability of that layer for new scenario data and previous scenario data is significantly different. In this case, we unfreeze that unstable layer to allow it to adapt to the new scenario (lines 26 to 27 in Algorithm 1), as the second unfreezing step in Figure 8b.
In the example in Figure 10, layers 1, 10, 15, and 40 are stable and can be frozen before the scenario change. After the scenario change, layers 1, 10, and 15 remain stable, while layer 40 becomes unstable. Consequently, we should resume training on layer 40 to let it quickly adapt to the new scenario.
4.2.3. Dynamic interval to conduct freezing processes.
SimFreeze continuously conducts freezing processes (i.e., tracks the CKA variation and freezes the converged layers). However, conducting the freezing process too frequently (e.g., every training iteration) can lead to excessive overhead, as CKA calculation requires time and energy. Therefore, we opt for periodically tracking the CKA variation of active layers and freezing the converging ones. Further, we propose using a logarithmic-based function that progressively decreases the interval before subsequent freezing process after each freezing process (line 10 in Algorithm 1). This approach takes into consideration that more and more layers converge as fine-tuning proceeds. For example, layer 10 in Figure 10 has not converged at the early training stage but gradually converged. The log-based function is calculated as $n=n*(1-1/log(n))$, where $n$ represents the interval to conduct a freezing process.
4.3. Utilizing Unlabeled Data
It is possible that a portion of the streaming-in training data arrives without labels, posing a challenge for traditional supervised fine-tuning processes. To address this issue and fine-tune the model in this case, EdgeOL adopts a semi-supervised learning technique (chapelle2009semi, ) to make use of both labeled and unlabeled data. In each fine-tuning round, EdgeOL first fine-tunes the model using unlabeled data via the self-supervised learning (chen2020simple, ) to improve the feature extraction ability of the model. Then, EdgeOL fine-tunes the model by supervised learning using the labeled data to improve the model’s performance in the target task (e.g., image classification for a particular dataset). This approach ensures EdgeOL’s robust performance and adaptability.
Figure 11. Overall fine-tuning execution time.